# SPAWNing Structural Priming Predictions from a Cognitively Motivated Parser

### v1

@@ 4 A case study: competing theories of reduced relative clauses
The Whiz-Deletion account of RCs (Chomsky, 1965) argues that the sub-tree corresponding to any RC, whether it is reduced or not, is headed by the same node: a complementizer phrase (CP). In full RCs, the lexical content in this phrase (the wh-word and auxiliary “was”) is overt, whereas in reduced RCs this lexical content is covert. The Participial-Phase account (Harwood, 2018) argues that full RCs are headed by CPs, and therefore include the wh-word and auxiliary “was”. Reduced passive and progressive RCs on the other hand are headed by Voice Phrase (VoiceP) and Progressive Phrase (ProgP) respectively. These nodes are also present in full RCs, but are nested deeper inside the CP. Consequently, the Participial-Phrase account, unlike the Whiz-Deletion account, does not assume the presence of a covert wh-word and auxiliary in reduced passive and progressive RCs. See § A.1 for trees that illustrate these differences.
@@
The differences between the two theories are implemented in the declarative memory (or grammar), summarized in § A.2; the procedural memory (parsing mechanism) is identical across both theories.
The differences in assumptions about the head of the RC are represented by different categories for nouns that are modified by RCs. In the Whiz-Deletion grammar, all nouns modified by RCs have the category NP/CP (i.e., a noun looking to combine with a CP on its right). In the Participial-Phrase grammar, nouns modified by FRCs have the NP/CP category, nouns modified by RRCs have the NP/VoiceP category, and nouns modified by ProgRRCs the NP/(VoiceP/ProgP) category.
The differences in assumptions about the presence of covert lexical items is represented by different null categories. For RCs which modify the subject of the clause, the Whiz-Deletion grammar has categories associated with the null wh-word, CP/(TPDP), and with the null auxiliary, (TPDP)/VoiceP, which when combined together result in the parse state CP/VoiceP, i.e., a state looking for a passive continuation. Both of these null categories are absent from the Participial-Phase grammar.444Both accounts have the same categories for the null wh-word in RCs which modify objects of clauses like “The cat the doctor examined was skittish”.We also experimented with a version of the Whiz-Deletion grammar that combines the two null categories into one, but found that this results in nearly identical predictions.
@@ 5 Methods
5.1 Priming Experimental paradigm
[…]
@@ 5.1 Priming Experimental paradigm
5.2 Estimating priming effects from humans
Participants
Design
@@ 5.2 Estimating priming effects from humans
Statistical model
We used regular expressions to classify all target completions into two categories: RRC vs. non-RRC. To estimate , we fit a Bayesian mixed-effects logistic regression model with the following three predictors (specified using Helmert contrasts) as fixed effects: All RCs vs. AMV, ProgRRC and FRC vs. RRC, and ProgRRC vs. FRC. We used a weakly informative prior and a maximal random effects structure with by-item and by-participant random intercepts and slopes (see Appendix D).
5.3 Generating priming predictions
We generated predictions from 12 types of models which varied along 3 dimensions: the grammar (Whiz-Deletion vs. Participial-Phase), amount of training data (0, 100, 500 sentences), and type of reanalysis (back to start vs. sample based on uncertainty). For each type, we created 1280 model instances by sampling different hyperparameters.
@@ 5.3 Generating priming predictions
There are two types of parameters in SPAWN. First, parameters that are fixed across all instances: decay ( in Equations 2, 5) set to , latency exponent ( in Equation 3) set to and maximum activation ( in Equation 4) set to . Second, parameters that differ for each model instance: latency factor ( in Equation 3) sampled from , and the noise parameter (the standard deviation of the normal distribution used to sample in Equation 1) sampled from . All of the parameters or the distributions they were sampled from, except for the noise parameter, were taken from Vasishth and Engelmann (2021). We sampled the noise parameter from instead of used in prior work because initial explorations revealed that when the noise parameter was sampled from this constrained range, models almost never retrieved syntactic categories with low base-level activation.
@@
To approximate the base-level and lexical activations that participants start with prior to the experiment, we trained SPAWN instances on either 100 or 500 sentences. These numbers are consistent with prior work which assume that participants start experiments with very weak priors (Delaney-Busch et al., 2019; Fine et al., 2010). We used templates to generate a dataset of 10000 sentences in which the relative frequency of different types of RC sentences mirrored corpus statistics from Roland et al. (2007); for example, only 1% of the training sentences contained an RRC (see Appendix § C) For each model instance, we sampled the training sentences from this dataset. Given the low probability of RRCs, many model instances never encountered these in their training data, and started with a base-level activation of for RRCs.
Priming experiment
We used the same stimuli lists from the human experiment to measure priming effects in the model instances, and the same statistical model to compute . On each prime trial, the model parsed the sentence and updated the base-level and lexical activations based on the final set of retrieved categories. On each target trial, the model parsed the partial prompt and we recorded the resulting parse state; the model was constrained to end with only one of two partial states: DP/PP (RRC parse) or TP/DP (active parse).555Other partial parse states would require completions with complex syntactic structures that people do not produce.
@@ 6 Results
Most participants (77%) never generated a single RRC target completion. Similarly most models (median of 66% across the 12 model types) never generated a single RRC parse state. Since the goal of this work is to find differences in the proportion of RRC completions across the primes, we only included in our statistical analyses the participants or models that generated at least one RRC completion or parse.
@@
The Participial-Phase model instances trained on no data with both re-analysis mechanisms captured the empirical priming pattern qualitatively; they also perfectly captured the magnitude of the priming effect in the RRC condition, and approximately captured the magnitude in the other conditions. With an increase in training data, the priming predictions diverged from the empirical data in two ways: First, the overall magnitude of priming decreased. Second, the relative differences between the priming conditions changed; this was especially salient in model instances which sampled the word to reanalyze based on uncertainty (as opposed to going back to the start) where the other RC conditions patterned more closely with the non-RC condition.
Whiz-Deletion vs. Humans
None of the Whiz-Deletion models captured the qualitative priming pattern observed in humans: the predicted effect for ProgRRC was higher than for FRC; empirically the priming effect in the two conditions was equal. The magnitude of priming in the RRC condition was also smaller than what was observed empirically.
@@ 7 Discussion
In this work we introduced a cognitively motivated parser, SPAWN, which can be used to generate quantitative behavioral predictions from any theory of syntax capable of generating incremental parse trees. SPAWN makes it possible to evaluate what theoretical differences (if any) result in differing sentence processing predictions. As a case study, we used SPAWN to generate predictions from two competing theories of reduced relative clauses (Whiz-Deletion and Participial-Phase) while modulating the reanalysis mechanism (going back to the start vs. sampling based on uncertainty) and the number of training examples (0, 100, 500). We compared the predictions from these different versions of the SPAWN model to human behavior from a large-scaled (N=769) web-based comprehension-to-production priming experiment.
We found that the predictions from the Whiz-Deletion SPAWN models did not capture the qualitative human priming behavior for any of the different reanalysis mechanisms or training sizes. In contrast, the Participial-Phase SPAWN models captured the human priming pattern, but only when the model was untrained (i.e., had no prior knowledge); As the prior knowledge increased, the relative differences between the priming conditions in humans and Participial-Phase models deviated, reinforcing assumptions from prior work that participants enter the experiment with weak prior knowledge. Taken together, these results suggest that the Participial-Phase account of reduced relative clauses captures the the structural representations people construct better than the Whiz-Deletion account.
@@ Conclusion
We proposed a framework to generate priming predictions from syntactic theory. Using relative clauses as a case study, we demonstrated how this framework can be used to adjudicate between competing syntactic theories and parsing mechanisms.
8 Limitations
Like all computational models, SPAWN makes many simplifying assumptions, such as assumptions about serial parsing or the assumptions underlying the specific mechanisms we implemented for reanalysis, “giving up” and prediction of null elements. While we explored the consequences of some of these assumptions when we modulated the reanalysis mechanism, this exploration was not exhaustive. It is important for future work to more systematically study the impact that the different simplifying assumptions have on behavioral predictions, which in turn can suggest concrete ways of improving the model. Since SPAWN is already capable of generating more fine grained predictions about the time take to read or reprocess speciic words in sentences — the time taken for any of the parsing steps (retrieval, integration and reanalysis) is measured in milliseconds by default in ACT-R — a promising direction for such a systematic study would be to evaluate these fine-grained predictions from SPAWN with empirical eye-tracking data. In a similar vein, applying this framework to other empirical phenomena can also shed light on how to improve the model.
Additionally, the current version of SPAWN is implemented using the CCG grammar formalism. Therefore, to use SPAWN to generate predictions from some theory, it is necessary to implement the structural assumptions in that theory using CCG notation. While CCG is a very flexible formalism that can be used to express the assumptions of most syntactic theories — for example, in this work we converted Minimalist Grammar trees to CCG trees — future versions of SPAWN can add support for other grammar formalisms to minimize the need for translating between grammar formalisms.
Appendix A Details about the two theories of reduced relative clauses
@@ A.1 Syntax trees
A.2 Differences in syntactic categories
@@
| “The cat being examined by … ” | NP/CP | NP/(VoiceP/ProgP) | |
| Null wh subject | “The cat NULLwh NULLaux examined by…” | CP/(TPDP) | MISSING |
@@
| Null progressive auxiliary | ‘The cat NULLwh NULLaux being examined by…” | (TPDP)/(VoiceP/ProgP) | MISSING |
@@ Appendix B SPAWN parsing details
B.1 Schematic of all the components
B.2 CCG combination and type-raising rules

### v2

@@ 4 A case study: Evaluating competing theories of reduced relative clauses
Under the Whiz-Deletion account of RCs (Chomsky, 1965), the sub-tree corresponding to any RC, whether reduced or not, is headed by the same node: a complementizer phrase (CP). In full RCs, the lexical content in this phrase (the wh-word and auxiliary “was”) is overt, whereas in reduced RCs this lexical content is covert. By contrast, under the Participial-Phase account (Harwood, 2018), while full RCs are headed by CPs, reduced passive and progressive RCs, are headed by Voice Phrase (VoiceP) and Progressive Phrase (ProgP) respectively. Consequently, the Participial-Phrase account, unlike the Whiz-Deletion account, does not assume the presence of a covert wh-word and auxiliary in reduced passive and progressive RCs. See § A.1 for trees that illustrate these differences.
@@
We implement two versions of SPAWN, a Whiz-Deletion version and a Participial-Phase version. The procedural memory (parsing mechanism) is identical across both versions. There are two main differences in the declarative memory (grammar) across the versions. First, they differ in the categories that nouns can be associated with: in the Whiz-Deletion version, all nouns modified by RCs have the category NP/CP (i.e., a noun looking to combine with a CP on its right), whereas in the Participial-Phase version, nouns modified by FRCs, RRCs and ProgRRCs are associated with different categories (NP/CP, NP/VoiceP and NP/ProgP respectively). Second, the versions have different null lexical items: the Whiz-Deletion version has lexical items for a null subject Wh-word, a null finite auxiliary and a null progressive auxiliary, all of which are absent in the Participial-Phase version (see § A.2).
@@ 5 Methods
5.1 Experimental paradigm
We used a comprehension-to-production priming paradigm to evaluate the two theories. In each experimental trial, human participants or SPAWN models were presented with three primes with the same structure, followed by an ambiguous partial prompt such as 5.1 that could be completed either with or without a reduced RC. We used four prime types: three prime types with RCs (one each for RRC, FRC and ProgRRC), as well as control primes without RCs, such as 5.1–5.1.
. The dog chased the boy and ran away.
@@ 5.1 Experimental paradigm
Estimating priming effects
We estimated priming effects by measuring the proportion of RRC target parses in the different priming conditions (see § 5.2 and § 5.3 for details on how these parses were measured in humans and models). Concretely, we estimated by fitting Bayesian mixed-effects logistic regression model with the following three predictors (specified using Helmert contrasts) as fixed effects: All RCs vs. AMV, ProgRRC and FRC vs. RRC, and ProgRRC vs. FRC. We used a weakly informative prior and a maximal random effects structure (see § E for further details).
@@
5.2 Experiment with human participants
Participants
Design
Measuring the proportion of RRC parses
We used regular expressions (§ F) to classify all target completions into two categories (RRC vs. non-RRC) and specified RRC completions as “success” in our Bayesian logistic regression model.
5.3 Experiment with SPAWN models
We generated predictions from 18 types of models which varied along 3 dimensions: the grammar (Whiz-Deletion vs. Participial-Phase; § 4, A.2), the reanalysis implementation (First-word regression and Entropy-Weighted reanalysis with temperature 1 or 10; § 3.2.3), and the number of training sentences (0, 100 or 1000 sentences). For each model type, we created 1280 model instances, which, as we describe below, share some hyperparameters and differ in others.
@@ 5.3 Experiment with SPAWN models
The following hyperparameters are fixed across all model instances: decay ( in Equations 2, 5), latency exponent ( in Equation 3), and maximum activation ( in Equation 4). The following hyperparameters differ for each model instance: latency factor ( in Equation 3), and the noise parameter ( in § 3.2.1).
The values for , , and as well as the sampling distributions for and were were taken from Vasishth and Engelmann (2021) (see § D for more details). We sampled from , because when sigma was sampled from Vasishth and Engelmann’s , some models never retrieved syntactic categories with low base-level activation. However, there were no qualitative differences in results between the two distributions (see § D)
@@
To set initial base-level and lexical activations of the models prior to the experiment, we trained the models on 0, 100, or 1000 sentences and updated the activations at the end of each sentence as described in § 3.3. These small numbers are consistent with prior work which assumes that participants start experiments with very weak priors (Delaney-Busch et al., 2019; Fine et al., 2010).
We used templates333https://github.com/grushaprasad/spawn/blob/main/create_training_dat.py to generate a dataset of 10000 sentences in which the relative frequency of different types of RC sentences mirrored corpus statistics from Roland et al. (2007); for example, only 1% of the training sentences contained an RRC (see § C for details about the distribution of sentence types). For each model instance, we sampled the training sentences from this dataset. Given the low probability of RRCs, many model instances never encountered these in their training data, and as such started with a base-level activation of for RRCs.
Design
We presented each model instance with the stimuli from the human experiment. On a prime trial, the model parsed the sentence and updated the base-level and lexical activations based on the final set of retrieved categories. On a target trial, the model parsed the partial prompt and we recorded the resulting parse state; the model was constrained to end with only one of two partial states: DP/PP (RRC parse) or TP/DP (active parse).
Measuring the proportion of RRC parses
We specified the DP/PP state as “success” in our Bayesian logistic regression model; unlike with humans, we do not need target completions to infer the parse the model assigned to the target.
@@ 6 Results
Most participants (77%) never generated a single RRC target completion. Similarly most models (median of 67% across the 18 model types) never generated a single RRC parse state. Since the goal of this work is to find differences in the proportion of RRC parses between the primes, we only included in our analyses and plots the participants or models that generated at least one RRC completion or parse state.
Human priming behavior
In the human experiment, we observed that the proportion of target RRC parses was highest when the target was preceded by RRC primes with the same structure, and lowest when preceded by AMV primes which did not have any relative clauses. The proportion of target RRC parses in other two priming conditions, ProgRRC and FRC, were equivalent relative to each other, lower than with RRC primes, and higher than with AMV primes. (Figure 4). See § E.3 for statistical analyses.
Whiz-Deletion vs. Humans
In the Whiz-Deletion models, processing ProgRRC sentences involves the retrieval of the same null complementizer as in RRC sentences, whereas processing FRC sentences does not (§ B.3). Consequently, these models predicted that the proportion of target RRC parses was greater with ProgRRC primes than with FRC primes (Figure 3), a pattern that does not align with the qualitative priming pattern observed in humans. Additionally, the magnitude of priming effects in the RRC condition were also generally smaller than what was observed in humans (Figure 3, bottom panel). These results together suggest that the Whiz-Deletion account of RRCs, at least the way we operationalized it, is not consistent with the representations humans build.
@@
In the Participial-Phase models, unlike in their Whiz-Deletion counterparts, processing ProgRRC, FRC, or AMV primes does not involve retrieving any categories that are shared with RRC sentences. However, the categories retrieved for ProgRRC and FRC but not AMV primes, increase the probability of reanalysis when processing the ambiguous target sentences (§ B.3). This reanalysis, as discussed in § 3.4, in turn increases the probability of the model eventually assigning an RRC parse to the target, especially if the models’ prior preference for AMV parses is relatively weak. Consequently, these models, particularly when they were trained on 0 and 100 sentences, predicted a graded effect which aligned with the qualitative priming pattern observed in humans: the proportion of target RRC parses was highest with RRC primes, followed by ProgRRC and FRC primes, and lowest with AMV primes. The models trained on sentences could not capture this qualitative pattern because they generated very few RRC sentences across the board. This suggests, in line with prior work (Delaney-Busch et al., 2019; Fine et al., 2010), that when modeling the production or processing of extremely infrequent structures (like RRCs), assuming weak prior knowledge might be necessary.
Of the models that captured the qualitative patterns, the models with first-word regression better captured the magnitude of the empirical priming effects (Figure 3). Taken together, these results suggest that, depending on the assumptions we make about reanalysis and strength of prior belief, the Participial-Phase account of RRCs, unlike the Whiz-Deletion account, can be consistent with the representations humans build.
@@ 7 Discussion
In this work we introduced a cognitively motivated parser, SPAWN, which can be used to generate quantitative behavioral predictions from contemporary syntactic theories that are based on lexicalized grammar formalisms. SPAWN makes it possible to evaluate what theoretical differences (if any) result in differing sentence processing predictions. As a case study, we used SPAWN to generate predictions from two competing theories of reduced relative clauses (Whiz-Deletion and Participial-Phase) while modulating the reanalysis mechanism and the number of training examples. We compared the predictions from these different versions of the SPAWN model to human behavior from a large-scale (N=769) web-based comprehension-to-production priming experiment.
We found that the predictions of the Whiz-Deletion SPAWN models did not capture the qualitative human priming behavior for any of the model types. In contrast, many of the Participial-Phase SPAWN models captured the qualitative patterns, with the models that best captured the magnitude of the empirical effects being ones with weak prior knowledge, that reprocesses the sentence from the beginning whenever reanalysis is triggered. Taken together, these results suggest that the Participial-Phase account of reduced relative clauses captures the structural representations people construct better than the Whiz-Deletion account.
Future work
This work tentatively suggests that first-word regression might better model human processing than entropy-weighted regression. This observation needs to be more robustly validated with other empirical phenomena (e.g., priming in PO/DO sentences). Additionally, some of the parsing mechanisms SPAWN implements, such as for reanalysis or predicting null elements, are likely too simplistic to account for human sentence processing more generally (see § G). Future work can tweak these mechanisms and evaluate the modified models against processing benchmarks like the SAP Benchmark Huang et al. (2024) which have more fine-grained measurements (e.g., reading time per word) across a range of psycholinguistic phenomena. Since the time taken for any of the parsing steps is measured in milliseconds by default in ACT-R, SPAWN can already generate quantitative predictions about the time taken to read or reprocess specific words in sentences, and therefore can be used with self-paced reading and eye-tracking datasets. Finally, future work can also use this paradigm to evaluate other competing syntactic theories.
@@ Conclusion
We proposed a cognitively plausible parser that can be used to generate quantitative behavioral predictions from syntactic theories. Using English reduced relative clauses as a case study, we demonstrated how this model can be used to adjudicate between competing syntactic theories and parsing mechanisms.
Acknowledgements
We would like to thank the anonymous reviewers, HSP 2023 and 2024 audience as well as Aniello De Santo, Shravan Vasishth, Will Merrill, Matt Wagers, Suhas Arehalli, Brian Dillon, Vijay Ramachandran and Joel Sommers for their valuable feedback.
This work was partly supported by an American Psychological Association Dissertation Research Award. The work was conducted using computational resources from the Maryland Advanced Research Computing Center (MARCC) and the Colgate Supercomputer (Partially funded by NSF Award #2346664).
Appendix A Details about the two theories of reduced relative clauses and how they are implemented in the declarative memory
@@ A.1 Syntax trees
A.2 Differences in syntactic categories between the two theories
@@
| “The cat being examined by … ” | NP/CP | NP/ProgP | |
| Null wh subject | “The cat NULLwh NULLpass examined by…” | CP/(TPDP) | MISSING |
@@
| Null progressive auxiliary | ‘The cat NULLwh NULLprog being examined by…” | (TPDP)/ProgP | MISSING |
When the noun in the Whiz-Deletion version combines with the null Wh subject and null finite or progressive auxiliary, it results in the same parse state as the Participial-Phase noun categories for RRC and ProgRRC: NP/VoiceP and NP/ProgP. We also explored an alternative implementation of the Whiz-Deletion account where instead of having three NULL categories — NULLWh, NULLpass and NULLprog — we had only two categories NULLWhpass (CP/VoiceP) and NULLWhprog (CP/ProgP). This implementation resulted in nearly identical results (§ 6)
Note, both accounts have the same categories for the null wh-word in RCs which modify objects of clauses like “The cat the doctor examined was skittish”: Null Complementizer (Object RCs) in the following table.
A.3 Syntactic categories shared by the two theories
| Category label | Example words | CCG rules |
| Determiner | the, a , an, some, his, her, many, a-lot-of | DP/NP |
| Determiner Phrase | something, everyone, non-violence, popularity | DP |
| Noun Phrase | dragon, media, palace, mission, trance, tax-fraud | NP |
| Preposition | on, to, into, by, at, in, down | PP/DP |
| Transitive verb (active) | accompanied, admired, betrayed, solved, forged | (TPDP)/DP |
| Transitive verb (passive) | accompanied, admired, betrayed, solved, forged | VoiceP/PP |
| Transitive verb (location object) | arrived, staggered, marched, participated | (TPDP)/PP |
| Intransitive verb | sang, cackled, complained, started-trending | TPDP |
| Complementizer (Subject RC) | who | CP/(TPDP) |
| Complementizer (Object RC) | who | CP/(((TPDP)/DP)/DP) |
| Null Complementizer (Object RC) | NULLWh | CP/(((TPDP)/DP)/DP) |
| Prog | being | ProgP/VoiceP |
| Auxiliary (followed by adjective) | was, were | (TPDP)/(NP/NP) |
| Auxiliary (finite) | was | (TPDP)/VoiceP |
| Auxiliary (progressive) | was | (TPDP)/ProgP |
| Adjective | unreliable, competent, well-known, signature, radical | NP/NP |
| Adverb | rapidly, diligently, in-surprise, sullenly, wistfully | TPTP |
| Conjunction | and | (TP/(TPDP))TP |
| EOS | . | endTP |
@@ Appendix B SPAWN parsing details
B.1 CCG combination and type-raising rules
@@
The parser starts by sequentially trying to apply each of the six composition rules, stopping once a successful combination is found. If no successful combination is found, then the parser tries to the type-raising rule and then sequentially apply all six composition rules.
B.2 Categories that can be followed by null elements
In the Whiz-Deletion grammar the NP/CP category, and the CP/(TPDP) can be followed by null elements, whereas in the Participial-Phase grammar, only the the NP/CP category can be followed by a null element (to account for object reduced RCs like “The cat the doctor examined was skittish”).
B.3 Analysis of example sentences with our grammar
| Old Parse state | Word | Correct category | Rule | New parse state |
| NULL | the | DP/NP | Initialize | DP/NP |
| DP/NP | cat | NP/VoiceP | Forward Harmonic Composition | DP/VoiceP |
| DP/VoiceP | examined | VoiceP/PP | Forward Harmonic Composition | DP/PP |
| DP/PP | by | PP/DP | Forward Harmonic Composition | DP/DP |
| DP/DP | the | DP/NP | Forward Harmonic Composition | DP/NP |
| DP/NP | doctor | NP | Forward Composition | DP |
| DP | liked | (TPDP)/DP | Type raise DP | |
| TP/(TPDP) | Forward Harmonic composition | TP/DP | ||
| TP/DP | the | DP/NP | Forward Harmonic composition | TP/NP |
| TP/NP | girl | NP | Forward Harmonic composition | TP |
| TP | EOS | endTP | Backward composition | end |
| Old Parse state | Word | Correct category | Rule | New parse state |
| NULL | the | DP/NP | Initialize | DP/NP |
| DP/NP | cat | NP/CP | Forward Harmonic Composition | DP/CP |
| DP/CP | NULLwh | CP/(TPDP) | Forward Harmonic Composition | DP/(TPDP) |
| DP/(TPDP) | NULLpass | (TPDP)/VoiceP | Forward Harmonic Composition | DP/VoiceP |
| DP/VoiceP | examined | VoiceP/PP | Forward Harmonic Composition | DP/PP |
| DP/PP | by | PP/DP | Forward Harmonic Composition | DP/DP |
| DP/DP | the | DP/NP | Forward Harmonic Composition | DP/NP |
| DP/NP | doctor | NP | Forward Composition | DP |
| DP | liked | (TPDP)/DP | Type raise DP | |
| TP/(TPDP) | Forward Harmonic composition | TP/DP | ||
| TP/DP | the | DP/NP | Forward Harmonic composition | TP/NP |
| TP/NP | girl | NP | Forward Harmonic composition | TP |
| TP | EOS | endTP | Backward composition | end |
| Old Parse state | Word | Correct category | Rule | New parse state |
| NULL | the | DP/NP | Initialize | DP/NP |
| DP/NP | cat | NP/ProgP | Forward Harmonic Composition | DP/ProgP |
| DP/ProgP | being | ProgP/VoiceP | Forward Harmonic Composition | DP/VoiceP |
| DP/VoiceP | examined | VoiceP/PP | Forward Harmonic Composition | DP/PP |
| DP/PP | by | PP/DP | Forward Harmonic Composition | DP/DP |
| DP/DP | the | DP/NP | Forward Harmonic Composition | DP/NP |
| DP/NP | doctor | NP | Forward Composition | DP |
| DP | liked | (TPDP)/DP | Type raise DP | |
| TP/(TPDP) | Forward Harmonic composition | TP/DP | ||
| TP/DP | the | DP/NP | Forward Harmonic composition | TP/NP |
| TP/NP | girl | NP | Forward Harmonic composition | TP |
| TP | EOS | endTP | Backward composition | end |
| Old Parse state | Word | Correct category | Rule | New parse state |
| NULL | the | DP/NP | Initialize | DP/NP |
| DP/NP | cat | NP/CP | Forward Harmonic Composition | DP/CP |
| DP/CP | NULLwh | CP/(TPDP) | Forward Harmonic Composition | DP/(TPDP) |
| DP/(TPDP) | NULLprog | (TPDP)/ProgP | Forward Harmonic Composition | DP/VoiceP |
| DP/ProgP | being | ProgP/VoiceP | Forward Harmonic Composition | DP/VoiceP |
| DP/VoiceP | examined | VoiceP/PP | Forward Harmonic Composition | DP/PP |
| DP/PP | by | PP/DP | Forward Harmonic Composition | DP/DP |
| DP/DP | the | DP/NP | Forward Harmonic Composition | DP/NP |
| DP/NP | doctor | NP | Forward Composition | DP |
| DP | liked | (TPDP)/DP | Type raise DP | |
| TP/(TPDP) | Forward Harmonic composition | TP/DP | ||
| TP/DP | the | DP/NP | Forward Harmonic composition | TP/NP |
| TP/NP | girl | NP | Forward Harmonic composition | TP |
| TP | EOS | endTP | Backward composition | end |
| Old Parse state | Word | Correct category | Rule | New parse state |
| NULL | the | DP/NP | Initialize | DP/NP |
| DP/NP | cat | NP/CP | Forward Harmonic Composition | DP/CP |
| DP/CP | which | CP/(TPDP) | Forward Harmonic Composition | DP/(TPDP) |
| DP/(TPDP) | was | (TPDP)/VoiceP | Forward Harmonic Composition | DP/VoiceP |
| DP/VoiceP | examined | VoiceP/PP | Forward Harmonic Composition | DP/PP |
| DP/PP | by | PP/DP | Forward Harmonic Composition | DP/DP |
| DP/DP | the | DP/NP | Forward Harmonic Composition | DP/NP |
| DP/NP | doctor | NP | Forward Composition | DP |
| DP | liked | (TPDP)/DP | Type raise DP | |
| TP/(TPDP) | Forward Harmonic composition | TP/DP | ||
| TP/DP | the | DP/NP | Forward Harmonic composition | TP/NP |
| TP/NP | girl | NP | Forward Harmonic composition | TP |
| TP | EOS | endTP | Backward composition | end |
| Old Parse state | Word | Correct category | Rule | New parse state |
| NULL | the | DP/NP | Initialize | DP/NP |
| DP/NP | cat | NP | Forward Composition | DP |
| DP | examined | (TPDP)/DP | Type raise DP | |
| TP/(TPDP) | Forward Harmonic Composition | TP/DP | ||
| TP/DP | the | DP/NP | Forward Harmonic Composition | TP/NP |
| TP/NP | doctor | NP | Forward Composition | TP |
| TP | and | (TP/(TPDP))TP) | Backward composition | TP/(TPDP) |
| (TP/(TPDP))TP) | liked | (TPDP)/DP | Forward Harmonic composition | TP/DP |
| TP/DP | the | DP/NP | Forward Harmonic composition | TP/NP |
| TP/NP | girl | NP | Forward Harmonic composition | TP |
| TP | EOS | endTP | Backward composition | end |